a particle having  a mass 0.5 kg is projected under gravity with a speed of 98 m/s at an angle of 60 degree. the magnitude of change of momentumin kgm/s of particle after 10 s is:
answer:49. but how?

Grade:12th pass

7 Answers

Samshad01
9 years ago
Samshad01
9 years ago
change in momentum = force × time
                                 = mg × time
                                 = 0.5 × 9.8 ×10 = 49 kgm/s      (g = 9.8m/s2)
shrilekha sharma
9 years ago
but the formula for change in momentum is given as 2musintheta in text? this formula holds true only when vertical motion is not involved that u have given but in  aprojectile both horizontal and vertical motions are involved ?? please explain?
Samshad01
9 years ago
Ok,
 Concider the horizontal motion of that object. There is no force is acting and momentum is conserved (I.e. change in momentum = 0).
On the vertical motion, the only force acting on the body is mg. Hence, according to second law of motion,
  Change in momentum = F*t
Change in momentum in verticle = 49 kgm/s.
Resultant2 = vx+ vy2
                    = 0 + 492
Resultant = 49 kgm/s
 
Samshad01
9 years ago
Ok,
 Concider the horizontal motion of that object. There is no force is acting and momentum is conserved (I.e. change in momentum = 0).
On the vertical motion, the only force acting on the body is mg. Hence, according to second law of motion,
  Change in momentum = F*t
Change in momentum in verticle = 49 kgm/s.
Resultant2 = vx+ vy2
                    = 0 + 492
Resultant = 49 kgm/s
 
Samshad01
9 years ago
Ok,
 Concider the horizontal motion of that object. There is no force is acting and momentum is conserved (I.e. change in momentum = 0).
On the vertical motion, the only force acting on the body is mg. Hence, according to second law of motion,
  Change in momentum = F*t
Change in momentum in verticle = 49 kgm/s.
Resultant2 = vx+ vy2
                    = 0 + 492
Resultant = 49 kgm/s
 
shrilekha sharma
9 years ago
thanks got it :) can u give me answer of my second question please?